Supplemental Operations Report June 16th

Supplemental Operations Report 16 June 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine/Astronomy Team visit

Non-nominal systems: None

Action taken for non-nominal systems: N/A

Generator: Offline and locked for the summer

ScienceDome Dual Split: On. AC at 76 degrees F to cool power system.

Solar— Operating 24/7

Solar— Nominal

Notes on power system: Balanced for summer.

Propane Reading, station tank – 76 %

Propane Reading, director tank— 64 %

Propane Reading, intern tank— 76 %

Propane Reading, generator— 78 %

Ethanol Free Gasoline – 0 gallons.

Water (Outpost tank) – 75 gallons

Notes on rovers: All rovers have been moved to town for the summer.

Notes on ATVs: All atvs have been stored onsite for summer

HabCar used and why, where? Yes, to town for water. Shifting column sometimes need to be pushed and held in an upward position in order for car to start. This has just happened and will need to be repaired before crew use in the fall. Also, the two passenger doors are not unlocking all the time.

CrewCar used and why, where? Yes, to Grand Junction to transport Martian Bio crew to return home. I am seeing increased instances when the power steering is not working when you first start moving. In addition, the fan in the heat/AC unit is extremely loud and part of the dash, including the radio has no power. The power switch for the glitchy window has broken and once I got the window to roll up, I left it that way for the summer. These things need to be repaired before crew use in the fall.

Luna used and why, where? No. Luna is not on campus until fall.

Campus wide inspection, if action taken, what and why? Everything is secure except the Hab which the Astronomy Team will secure when they leave.

General notes and comments: Summer is hot!

Summary of general operations: We had a school visit on Tuesday and the astronomy team arrived on Wednesday and leaves on Saturday. This will be the last of our field season’s official schedule.

Summary of internet: Starlink currently operational at the Hab. The Astronomy Team will unplug it and move the dish inside for the summer.

Summary of suits and radios: Radios stored for the summer. Batteries are in the ScienceDome. Spacesuits with NorCal Chapter for summer.

Summary of Hab operations: All food has been removed for the summer. Rodenticide to be placed on upper and lower decks when the Astronomy Team leaves. They will be at the Hab until Saturday and I have given them a list of the remaining things they need to do to shut down the Hab when they leave. A new lock will be installed.

Summary of Outpost operations: Intern trailer windows have been covered inside with reflective bubble wrap. Rodenticide placed under kitchen sink and in storage compartment. Locked. My trailer will be shut down and locked when I leave. Rodenticide has been placed under my kitchen counter. Both toilet tanks emptied.

Summary of GreenHab operations: Shutdown for the summer. No rodenticide is in this building.

Summary of ScienceDome operations: AC on to 76 degrees F. Window covered and building shut down for the summer. New lock installed. Rodenticide placed under the sink just in case any rodents are in there.

Summary of RAM operations: Shut down for summer. New lock installed. All power has been turned off. Rodenticide placed in an open plastic container by the roll up door.

Summary of any observatory issues: Operational once again!

Summary of health and safety issues: Rodenticide has been placed in all buildings for the summer. In most cases it is safe from domestic animals, but should be checked and any remaining material removed at the beginning of the field season in October.

Copyright © The Mars Society. All rights reserved. | Main Site